# Open-Set Video-based Facial Expression Recognition with Human Expression-sensitive Prompting

### v1

@@ Abstract.
In Video-based Facial Expression Recognition (V-FER), models are typically trained on closed-set datasets with a fixed number of known classes. However, these V-FER models cannot deal with unknown classes that are prevalent in real-world scenarios. In this paper, we introduce a challenging Open-set Video-based Facial Expression Recognition (OV-FER) task, aiming at identifying not only known classes but also new, unknown human facial expressions not encountered during training. While existing approaches address open-set recognition by leveraging large-scale vision-language models like CLIP to identify unseen classes, we argue that these methods may not adequately capture the nuanced and subtle human expression patterns required by the OV-FER task. To address this limitation, we propose a novel Human Expression-Sensitive Prompting (HESP) mechanism to significantly enhance CLIP’s ability to model video-based facial expression details effectively, thereby presenting a new CLIP-based OV-FER approach. Our proposed HESP comprises three components: 1) a textual prompting module with learnable prompt representations to complement the original CLIP textual prompts and enhance the textual representations of both known and unknown emotions, 2) a visual prompting module that encodes temporal emotional information from video frames using expression-sensitive attention, equipping CLIP with a new visual modeling ability to extract emotion-rich information, and 3) a delicately designed open-set multi-task learning scheme that facilitates prompt learning and encourages interactions between the textual and visual prompting modules, augmenting the understanding of novel human emotion-related patterns in video sequences to a great extent. Extensive experiments conducted on four OV-FER task settings demonstrate that HESP can significantly boost CLIP’s performance (a relative improvement of 17.93% on AUROC and 106.18% on OSCR) and outperform other state-of-the-art open-set video understanding methods by a large margin.
@@
ccs: Human-centered computingccs: Computing methodologies Computer vision
@@ 1. Introduction
Video-based Facial Expression Recognition (V-FER) targets the recognition of predetermined emotion categories, such as anger, happiness, and sadness, from video data. Despite its advancements in enhancing human-centered visual comprehension, this method often operates within a confined set of predefined emotion categories. However, this limitation fails to fully encapsulate the intricate and diverse spectrum of human expressions, thereby constraining the potential to encompass a broader range of emotional states. To address this limitation, we introduce the Open-set Video-based Facial Expression Recognition (OV-FER) task, which seeks to identify known and discover new human emotions from videos. Similar to other open-set paradigms (Kong and Ramanan, 2021; Chen et al., 2021; Bendale and Boult, 2016; Moon et al., 2022; Neal et al., 2018), the OV-FER requires the ability to recognize and classify previously unseen or ‘unknown’ emotional expressions. Rooted in human emotion understanding, an effective OV-FER model is supposed to provide a robust understanding of various emotion and facial expression patterns in diverse and unpredictable environments, thus benefiting diversified applications like smart healthcare (Bisogni et al., 2022), human-computer interaction (Chattopadhyay et al., 2020), and driver assistance system (Wilhelm, 2019).
@@
Recently, several studies (Wang et al., 2023; Ren et al., 2023; Liao et al., 2022) have proposed leveraging large-scale pre-trained models like CLIP (Radford et al., 2021) to facilitate open-set recognition tasks. With CLIP’s remarkable generalization capabilities, it becomes feasible to effectively identify new categories from open-set data using appropriate text prompts. Consequently, CLIP has emerged as a valuable tool for open-set analysis, demonstrating its efficacy in handling unseen data instances (Wang et al., 2023; Ren et al., 2023; Liao et al., 2022). However, when applying CLIP to video analysis, a critical issue arises: CLIP was originally designed for image-text associations, and its support for video analysis, particularly in the context of human emotion videos, is not addressed appropriately in the vanilla model. To tackle this, a few studies (Zara et al., 2023; Wu et al., 2024) attempt to capture spatial-temporal dynamic information in videos to extend the capability of CLIP on open-set video understanding tasks. For instance, AutoLabel (Zara et al., 2023) introduces an additional self-attention temporal pooling for CLIP to aggregate frame-level features as output video-level features. Open-VCLIP (Wu et al., 2024) proposes a temporal attention view of each self-attention layer in CLIP to facilitate the aggregation of global temporal information for open-set video action recognition. Despite promising progress, these CLIP video-based methods still achieve unsatisfactory results on OV-FER, where changes in facial expressions both between different videos and within videos can be very subtle (Liu et al., 2023), as shown in Fig. 1.
@@
To address the above problems when applying the CLIP on OV-FER, we propose a novel human expression-sensitive prompting (HESP) mechanism to significantly upgrade the CLIP by equipping it with the ability to understand facial details and model subtle human expressions effectively, therefore enabling the CLIP-based model to properly deal with unseen emotion categories in the OV-FER task.
In general, our contributions are summarized as follows:
@@
  • We are the first research to explore the OV-FER task. To tackle the task, we propose a novel Human Expression-sensitive Prompting (HESP) framework comprising textual and visual prompting modules, along with a multi-task learning component. HESP significantly augments CLIP’s efficacy in achieving effective OV-FER and, to our knowledge, is the first CLIP-based approach for the OV-FER task.
@@
  • Methodologically, first, in our textual prompting, we introduce learnable prompt representations to help better depict human emotions in text inputs for both known classes in closed-set datasets and the novel, unknown class in open-set datasets, meanwhile, for visual prompting, we also devise a novel, learnable visual prompting to help CLIP attend to the expression-sensitive areas on different frames of the video, providing temporal visual expression-sensitive information for both known and unknown data. Then, an open-set multi-task learning scheme is proposed to facilitate the cross-modal prompt interaction, thus further enhancing the HESP’s performance in capturing inter-video emotion differences for both known and unknown categories.
@@ 2. Related work
Open-set Recognition. Current open-set recognition methods are classified as: discriminative models and generative models. Discriminative models construct auxiliary blocks and add rules to reject unknown classes (Bendale and Boult, 2016; Chen et al., 2021; Huang et al., 2022). Generative models generate fake data that simulates novel classes during training (Kong and Ramanan, 2021; Moon et al., 2022). Although encouraging results have been achieved, both types of methods show limited generalization abilities to various downstream tasks like FER due to the subtle inter-class differences in facial expressions. Recently, there have been studies (Zhu et al., 2023; Zhang et al., 2024) addressing the issue of open-set image FER. For example, VBExT (Zhu et al., 2023) designs a Bi-Attention Expression Transformer to extract more compact features. Open-set FER (Zhang et al., 2024) treats it as a noisy label detection problem and utilizes attention map consistency to separate open-set samples. However, these methods do not have the ability to handle videos, which affects their applications to OV-FER.
Large-scale Vision-language Models and Prompting Learning. Pre-trained large-scale vision-language models have become significant in computer vision (CV). These models, for instance CLIP (Radford et al., 2021) and GPT-4 (Achiam et al., 2023), learning rich visual and linguistic features through pre-training on large-scale image and text data, achieving significant performance improvements in various downstream tasks, such as image classification (Zhou et al., 2022), object detection (Vidit et al., 2023), as well as open-set recognition (Wang et al., 2023; Ren et al., 2023; Liao et al., 2022). Prompt learning (PL) aids large-scale models in adapting to various downstream tasks (Zhou et al., 2022; Tsimpoukelli et al., 2021; Yao et al., 2024), which is classified into fixed textual prompts, learnable textual prompts, learnable visual prompts, and learnable vector sequence visual prompts. CoOp (Zhou et al., 2022) employs learnable text prompts, avoiding manual prompt tuning. VP (Bahng et al., 2022) uses pixel-level learnable perturbation visual prompts, proving the effectiveness of perturbation for CLIP. VPT (Jia et al., 2022) uses trainable parameters as vector sequence visual prompts to improve the performance. Despite the progress, existing PL methods are mainly designed for generic image recognition, lacking effective PL for more challenging FER tasks.
@@ 3.1. Preliminary
Negative representations, opposite to known class representations, are particularly beneficial for discovering unknown patterns in open-set datasets (Chen et al., 2021; Feng et al., 2023; Chen et al., 2020). In open-set data, negative representations help distinguish irrelevant patterns, like background features, enabling better recognition of meaningful, unknown patterns. Therefore, data with patterns can be collected from negative representations of all categories to discover unknown categories. For more details, we refer readers to (Chen et al., 2021). In the OV-FER task, since the differences between videos could be very small, it is important to identify and de-emphasize the patterns related to negative representations, thereby driving us to follow such a mechanism for implementing the CLIP-based OV-FER approach.
@@
where $F_{T}$ and $F_{V,i}$ are extracted textual and visual features, respectively. Then, CLIP correlates the text features with the visual features based on the relationships between the input texts and images. The correlation between $F_{T}$ and $F_{V,i}$ is estimated by $corr(F_{T},F_{V,i})$ where $corr$ is a correlation estimation process. Normally, $corr$ is implemented by inner-product followed by a softmax, and the obtained scores can be used for various zero-shot purposes, including open-set tasks. For example, if the correlation score between an image and the phrase ”a photo of a happy expression” is very high, we can consider that this image contains a happy face, even without training or fine-tuning.
@@ 3.3. Textual Prompting
To discover unknown classes in open-set data, we follow (Chen et al., 2021) and introduce negative representations. Specifically, we introduce $K$ different tokens, each of which helps discover patterns that do NOT belong to a specific known emotion category. In practice, we apply $K$ fixed negative textual prompts as the tokens to help explore unknown emotion categories in open-set datasets. The negative textual prompts $\bar{T}(k)$ are usually in the form of ”This video is not [CLASS-k]”, where the [CLASS-k] enumerates all the closed-set emotional categories. Then, we employ the the CLIP text encoder $\phi_{T}$ to obtain the negative textual features $\bar{F}_{T^{\prime}}=\{\phi_{T}(\bar{T}(1)),\ldots,\phi_{T}(\bar{T}(K))\}$. Since the negative textual representations mainly summarize irrelevant patterns, we simply use fixed templates. We can demonstrate in the experiment that utilizing fixed negative texts can already obtain good performance.
@@ 3.4. Visual Prompting
For open-set data, we still introduce negative representations. Instead of textual prompting where we have an explicit negation phrase ”not”, we attempt to apply implicit negative visual representations $\bar{V}_{i,k}$ that need to be optimized together with our HESP promoting modules. Specifically, given $K$ known categories, we start with randomly initialized $K$ tensors to represent the negative representations for each $K$ known category. To ensure consistency with closed-set data inputs for CLIP, we set the dimensions of each implicit negative visual representation as the same with a frame $V_{i}(q)$, namely $3\times 224\times 224$. Then, using the encoder $\phi_{V}$ in CLIP, we extract the negative visual prompting features as $\bar{F}_{V}^{\prime}=\{\phi_{V}(\bar{V}_{i,1}),\ldots,\phi_{V}(\bar{V}_{i,K})\}$. After training, although the negative text representations already facilitate the identification of irrelevant visual patterns, the performance could be affected by visual noises. Therefore, further calculating the distances between open-set video input and learned implicit negative visual representations allows us to better identify irrelevant visual patterns and help discover more appropriate unknown patterns. We can show in supplementary to support this design.
@@ 4.1.1. Datasets
AFEW (Dhall et al., 2015): The AFEW dataset has served as the evaluation platform for EmotiW since 2013. It contains videos from movies and TV series with spontaneous expressions under various conditions. With seven basic emotion labels, i.e., anger, disgust, fear, happiness, sadness, surprise, and neutral, AFEW comprises Train (738 videos), Val (352 videos), and Test (653 videos) splits. Due to the absence of test labels, we followed current methods (Liu et al., 2022b; Zhao and Liu, 2021), using only the Training/Val set for evaluation.
@@ 4.1.3. Evaluation Protocols
Following the existing work (Chen et al., 2021; Zhu et al., 2023), we selected the area under the receiver operating characteristic curve (AUROC) and the open-set classification rate (OSCR) as our evaluation metrics. While AUROC solely assesses the model’s capacity to differentiate between known and unknown classes, OSCR not only gauges this distinction but also evaluates the accuracy of known class recognition. Higher values indicate better model performance.
@@ 4.4. Visualization
4.4.1. Visualization of Prediction Results.
In order to more intuitively demonstrate the accuracy of our method, we visualized the model prediction results and compared the CLIP+ARPL, Open-VCLIP (Wu et al., 2024), and our HESP in Fig. 4. The prediction results show that our HESP can not only accurately identify known classes, but also detect unknown emotion classes, effectively.
Figure 4. Prediction results of different methods.

### v2

@@ Abstract.
In Video-based Facial Expression Recognition (V-FER), models are typically trained on closed-set datasets with a fixed number of known classes. However, these models struggle with unknown classes common in real-world scenarios. In this paper, we introduce a challenging Open-set Video-based Facial Expression Recognition (OV-FER) task, aiming to identify both known and new, unseen facial expressions. While existing approaches use large-scale vision-language models like CLIP to identify unseen classes, we argue that these methods may not adequately capture the subtle human expressions needed for OV-FER. To address this limitation, we propose a novel Human Expression-Sensitive Prompting (HESP) mechanism to significantly enhance CLIP’s ability to model video-based facial expression details effectively. Our proposed HESP comprises three components: 1) a textual prompting module with learnable prompts to enhance CLIP’s textual representation of both known and unknown emotions, 2) a visual prompting module that encodes temporal emotional information from video frames using expression-sensitive attention, equipping CLIP with a new visual modeling ability to extract emotion-rich information, and 3) an open-set multi-task learning scheme that promotes interaction between the textual and visual modules, improving the understanding of novel human emotions in video sequences. Extensive experiments conducted on four OV-FER task settings demonstrate that HESP can significantly boost CLIP’s performance (a relative improvement of 17.93% on AUROC and 106.18% on OSCR) and outperform other state-of-the-art open-set video understanding methods by a large margin. Code is available at https://github.com/cosinehuang/HESP.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681583isbn: 979-8-4007-0686-8/24/10ccs: Human-centered computingccs: Computing methodologies Computer vision
@@ 1. Introduction
Video-based Facial Expression Recognition (V-FER) targets the recognition of predetermined emotion categories from video data. Despite its advancements in enhancing human-centered visual comprehension, this method usually focuses on a limited set of predefined categories. This restriction does not capture the full complexity and diversity of human expressions, limiting its ability to recognize a broader range of emotional states. To address this limitation, we introduce the Open-set Video-based Facial Expression Recognition (OV-FER) task, which seeks to identify known and discover new human emotions from videos. Similar to other open-set paradigms (Kong and Ramanan, 2021; Chen et al., 2021; Bendale and Boult, 2016; Moon et al., 2022; Neal et al., 2018), the OV-FER requires the ability to recognize and classify previously unseen or ‘unknown’ expressions. Rooted in human emotion understanding, an effective OV-FER model is supposed to provide a robust understanding of various facial expression patterns in diverse and unpredictable environments, thus benefiting diversified applications like smart healthcare (Bisogni et al., 2022), human-computer interaction (Chattopadhyay et al., 2020), and driver assistance system (Wilhelm, 2019).
@@
Recently, several studies (Wang et al., 2023; Ren et al., 2023; Liao et al., 2022) have proposed leveraging large-scale pre-trained models like CLIP (Radford et al., 2021) to facilitate open-set recognition tasks. With CLIP’s remarkable generalization capabilities, it becomes feasible to effectively identify new categories from open-set data. However, when applying CLIP to video analysis, a critical issue arises: CLIP was originally designed for image-text associations, and its support for video analysis, particularly in the context of human emotion videos, is not addressed appropriately in the vanilla model. To tackle this, a few studies (Zara et al., 2023; Wu et al., 2024) attempt to capture spatial-temporal dynamic information in videos to extend the capability of CLIP on open-set video understanding tasks. For instance, AutoLabel (Zara et al., 2023) introduces an additional self-attention temporal pooling for CLIP to aggregate frame-level features as video-level features. Open-VCLIP (Wu et al., 2024) proposes a temporal attention view of each self-attention layer in CLIP to facilitate the aggregation of global temporal information. Despite promising progress, these CLIP video-based methods still achieve unsatisfactory results on OV-FER, where changes in facial expressions both between and within videos can be very subtle (Liu et al., 2023), as shown in Fig. 1.
@@
To address the above problems when applying the CLIP on OV-FER, we propose a novel human expression-sensitive prompting (HESP) mechanism to significantly upgrade the CLIP by equipping it with the ability to understand facial details and model subtle human expressions effectively, therefore enabling the CLIP-based model to properly deal with unseen emotion categories in the OV-FER task. In general, our contributions are summarized as follows:
@@
  • We are the first research to explore the OV-FER task. To tackle the task, we propose a novel Human Expression-sensitive Prompting (HESP) framework comprising textual and visual prompting modules, along with a multi-task learning component. HESP significantly augments CLIP’s efficacy in achieving effective OV-FER.
@@
  • Methodologically, first, we introduce learnable text prompts to better depict emotions in both known and unknown classes. We also create learnable visual prompts to guide CLIP in focusing on expression-sensitive areas in video frames, capturing temporal visual expression-sensitive information for both known and unknown data. Then, an open-set multi-task learning scheme is proposed to facilitate the cross-modal prompt interaction, thus further enhancing the HESP’s performance in capturing inter-video emotion differences for both known and unknown categories.
@@ 2. Related work
Open-set Recognition. Current open-set recognition methods are classified as: discriminative models and generative models. Discriminative models adds rejection rules to identify unknown classes (Bendale and Boult, 2016; Chen et al., 2021; Huang et al., 2022). Generative models generate fake data that simulates novel classes during training (Kong and Ramanan, 2021; Moon et al., 2022). Although encouraging results have been achieved, both types of methods show limited generalization abilities to FER task due to the subtle inter-class differences in facial expressions. Recently, there have been studies (Zhu et al., 2023; Zhang et al., 2024) addressing the issue of open-set image FER. For example, VBExT (Zhu et al., 2023) designs a Bi-Attention Expression Transformer to extract more compact features. Open-set FER (Zhang et al., 2024) treats it as a noisy label detection problem and utilizes attention map consistency to separate open-set samples. However, these methods are not designed to handle videos, which limits their applicability in OV-FER.
Large-scale Vision-language Models and Prompting Learning. Pre-trained large-scale vision-language models have become significant in computer vision (CV). These models, for instance CLIP (Radford et al., 2021) and GPT-4 (Achiam et al., 2023), learning rich visual and linguistic features through pre-training on large-scale image and text data, achieving significant performance improvements in various downstream tasks, such as image classification (Zhou et al., 2022), object detection (Vidit et al., 2023), as well as open-set recognition (Wang et al., 2023; Ren et al., 2023; Liao et al., 2022). Prompt learning (PL) aids large-scale models in adapting to various downstream tasks (Zhou et al., 2022; Tsimpoukelli et al., 2021; Yao et al., 2024). CoOp (Zhou et al., 2022) employs learnable text prompts, avoiding manual prompt tuning. VPT (Jia et al., 2022) uses trainable parameters as vector sequence visual prompts to improve the performance. Despite the progress, existing PL methods are mainly designed for generic image recognition, lacking effective PL for more challenging FER tasks.
@@ 3.1. Preliminary
Negative representations, opposite to known class representations, are particularly beneficial for discovering unknown patterns in open-set datasets (Chen et al., 2021; Feng et al., 2023; Chen et al., 2020). In open-set data, negative representations help distinguish irrelevant patterns, like background features, enabling better recognition of meaningful, unknown patterns. Therefore, data with patterns can be collected from negative representations of all categories to discover unknown categories. For more details, we refer readers to (Chen et al., 2021). In the OV-FER task, since the differences between videos could be very small, it is important to identify the patterns related to negative representations, thereby driving us to follow such a mechanism for implementing the CLIP-based OV-FER approach.
@@
where $F_{T}$ and $F_{V,i}$ are extracted textual and visual features, respectively. Then, CLIP correlates the text features with the visual features based on the relationships between the input texts and images. The correlation between $F_{T}$ and $F_{V,i}$ is estimated by $corr(F_{T},F_{V,i})$ where $corr$ is a correlation estimation process. Normally, $corr$ is implemented by inner-product followed by a softmax, and the obtained scores can be used for various zero-shot purposes, like open-set tasks. For example, if the correlation score between an image and the phrase ”a photo of a happy expression” is very high, we can consider that this image contains a happy face, even without training or fine-tuning.
@@ 3.3. Textual Prompting
To discover unknown classes in open-set data, we follow (Chen et al., 2021) and introduce negative representations. Specifically, we introduce $K$ different tokens, each of which helps discover patterns that do NOT belong to a specific known emotion category. In practice, we apply $K$ fixed negative textual prompts as the tokens to help explore unknown emotion categories in open-set datasets. The negative textual prompts $\bar{T}(k)$ are usually in the form of ”This video is not [CLASS-k]”, where the [CLASS-k] enumerates all the closed-set emotional categories. Then, we employ the the CLIP text encoder $\phi_{T}$ to obtain the negative textual features $\bar{F}_{T^{\prime}}=\{\phi_{T}(\bar{T}(1)),\ldots,\phi_{T}(\bar{T}(K))\}$.
@@ 3.4. Visual Prompting
For open-set data, we still introduce negative representations. Instead of textual prompting where we have an explicit negation phrase ”not”, we attempt to apply implicit negative visual representations $\bar{V}_{i,k}$ that need to be optimized together with our HESP promoting modules. Specifically, given $K$ known categories, we start with randomly initialized $K$ tensors to represent the negative representations for each $K$ known category. To ensure consistency with closed-set data inputs for CLIP, we set the dimensions of each implicit negative visual representation as the same with a frame $V_{i}(q)$, namely $3\times 224\times 224$. Then, using the encoder $\phi_{V}$ in CLIP, we extract the negative visual prompting features as $\bar{F}_{V}^{\prime}=\{\phi_{V}(\bar{V}_{i,1}),\ldots,\phi_{V}(\bar{V}_{i,K})\}$. After training, although the negative text representations already facilitate the identification of irrelevant visual patterns, the performance could be affected by visual noises. Therefore, further calculating the distances between open-set video input and learned implicit negative visual representations allows us to better identify irrelevant visual patterns and help discover more appropriate unknown patterns.
@@ 4.1.1. Datasets
AFEW (Dhall et al., 2015): The AFEW dataset contains videos from movies and TV series with spontaneous expressions under various conditions. With seven basic emotion labels, i.e., anger, disgust, fear, happiness, sadness, surprise, and neutral, AFEW comprises Train (738 videos), Val (352 videos), and Test (653 videos) splits.
@@ 4.1.3. Evaluation Protocols
Following the existing work (Chen et al., 2021; Zhu et al., 2023), we selected the Area Under the Receiver Operating Characteristic Curve (AUROC) and the Open-Set Classification Rate (OSCR) as our evaluation metrics.
@@ 4.3. Ablation Studies
4.3.4. Experiments on Large Dataset and Inference Speed.
To further validate the effectiveness and universality of our method, we evaluated our HESP on a large dataset DFEW (Jiang et al., 2020) containing 7 basic emotions, and the results are shown in Table 7. Our method still performs best, relatively surpassing Open-VCLIP (Wu et al., 2024) 17.63% and 36.30% on AUROC and OSCR, respectively. Meanwhile, we compared the full model inference speed of various methods in Table 7. Analysis of the table reveals that our method significantly improves model performance while reducing time consumption.
Table 7. OV-FER evaluation on the DFEW dataset.
Method Performance Computation cost
AUROC OSCR Inference speed (s)
ARPL (Chen et al., 2021) 54.46 29.12 0.021
Open-set FER (Zhang et al., 2024) 54.78 27.78 0.077
Open-VCLIP (Wu et al., 2024) 56.83 39.92 0.131
Baseline 2+HESP (ours) 66.85 54.41 0.028
@@ 5. Conclusion
Acknowledgements.
This work was supported by the National Natural Science Foundation of China grant (62076227), Natural Science Foundation of Hubei Province grant (2023AFB572) and Hubei Key Laboratory of Intelligent Geo-Information Processing (KLIGIP-2022-B10).